Two pieces of copper with equal mass are placed in a well-insulated container of negligible heat
capacity and allowed to come to equilibrium. If one of the pieces had an initial temperature of
144 °C and the other’s initial temperature was 72.0°C, what will the final temperature in the
container be? (cCu(s) = 0.385 J⁄g ∙ K) State any assumptions needed to solve the problem.
Answer #1

Q = mc∆T

Q = heat energy (Joules, J), m = mass of a substance (g)

c = specific heat (units J/g∙oC), is a symbol meaning "the change in"

∆T = change in temperature (oC Celcius)

Let us take each 1 gm copper

specific heat capacity of copper = 0.385 J⁄g ∙ K

∆T = Tf- Ti

Tf = final temperature Ti = initial temperature

Heat gained by low temperature copper = Heat lost by high temperature copper

1 x 0.385 X (144 - Tf) = 1 x 0.385 x (Tf -72)

144 + 72 = 2 Tf

Tf = 108oC

Hence  the final temperature in the container will be 108oC